OpenStudy (igster1112):

Solve the equation -(x - 10)=7

OpenStudy (zkoqmosii):

Solve for x: 10-x = 7 Isolate terms with x to the left hand side. Subtract 10 from both sides: (10-10)-x = 7-10 Look for two terms that sum to zero. 10-10 = 0: -x = 7-10 Evaluate 7-10. 7-10 = -3: -x = -3 Multiply both sides by a constant to simplify the equation. Multiply both sides of -x = -3 by -1: -x/(-1) = 3 Any non zero number divided by itself is one. (-1)/(-1) = 1: Answer: x = 3